THE MICROWAVE SPECTRUM AND DIPOLE MOMENT OF CIS-HEX-3-ENE-1,5-DIYNE

\underline{ROBB J. WILSON}, ROBERT L. KUCZKOWSKI, Department of Chemistry, University of Michigan, Ann Arbor, Michigan 48109-1055, U.S.A.; RYAN L. FIMMEN, ROBERT J. McMAHON, Department of Chemistry, University of Wisconsin, Madison, Wisconsin 53706-1396, U.S.A..

The microwave spectrum of cis-hex-3-ene-1,5-diyne has been observed using a Fourier transform spectrometer. Twentytwo transitions were fit to eight constants to 3.9 kHz. The rotational constants have been determined as A = 6955.300(2) MHz, B = 2621.515(1) MHz, and C = 1900.734(1) MHz. The constants were calculated using ab initio methods at the MP2/6-31G* level as A = 6906.098 MHz, B = 2579.524 MHz, and C = 1878.047 MHz showing reasonable agreement with experiment. The electric dipole moment was measured to be 0.18(1) D. This molecule is the parent molecule of a series of compounds which undergo the ’Bergman cyclization’ reaction, and thus serves as a precursor to aromatic compounds. It also has important implications in biological systems, and may exist in the interstellar medium.